Question:

If π‘Ž and 𝑏 are two sides of a right angled hypotenuse 𝑐 and let 𝑝 be the perpendicular from the opposite

vertex on the hypotenuse, show that:

1. (𝛿 𝑝/π›Ώπ‘Ž)bΒ =b3/c3

2. (𝛿𝑝/π›Ώπ‘Ž )A=b/c



1
Expert's answer
2021-08-17T09:39:20-0400

Part 1

From Pythagoras theorem

"c^2=a^2+b^2\\"

From the similarity of triangles

"\frac{p}{a}=\frac{b}{c}"

When we differentiate partially with respect to a

"\frac{\partial p}{\partial a}=\frac{b}{c}"

Cubing each side of the equation and neglecting the smallest differentials we get

"\frac{\partial p}{\partial a}=(\frac{b}{c})^3\\
\frac{\partial p}{\partial a}=\frac{b^3}{c^3}"


Part 2

From Pythagoras theorem

"c^2=a^2+b^2\\"

From the similarity of triangles

"\frac{p}{a}=\frac{b}{c}"

When we differentiate partially with respect to a

"\frac{\partial p}{\partial a}=\frac{b}{c}"